# NCoder - A Quantum Field Theory approach to encoding data

### v1

@@ 1 Introduction
To address our various questions we introduce in this note the NCoder – a neural network built around our aim for interpretability. The NCoder can be viewed as a deconstructed autoencoder in which the ‘latent layer’ is not built during the simultaneous optimization of the encoder/decoder but rather is fixed before training. Although this constraint limits the freedom of the network, it improves its ‘readability’. Motivated by the central role played by $n$-point functions in the analysis of field theories and general data science models, the NCoder uses the space of correlation functions as its prescribed latent layer. This set up renders the NCoder a potentially powerful tool for analyzing the importance of interpretable statistics in data generation and classification tasks. The paper is organized as follows: in Section 2 we review relevant concepts from probabiltiy and statistics such as the definition of moments and cumulants. In Section 2.2 we introduce the moment/cumulant reconstruction problem, and present a general solution in terms of the Edgeworth expansion which generalizes ideas from perturbation theory to arbitrary probability models. In Section 3 we briefly review how image data can be regarded as draws from a lattice field theory in preparation for the analysis of the latter sections. Section 4 presents the NCoder in detail, with a special focus on rendering the task of obtaining $n$-point functions feasible in light of the computational cost of the combinatorics involved. In Section 5 we use the NCoder to perform image generation tasks involving the MNIST and Fashion MNIST data sets. We conclude in Section 6 in which we discuss our findings, and suggest interesting directions for future work.
@@ 2.2 The reconstruction problem and the Edgeworth expansion
In what follows it we will often work under the assumption that $p_{0}$ is a Gaussian distribution whose first and second cumulants match with those of $p$. In this case (20) takes the form
@@ 4.1 Review of autoencoders
In most cases $m_{\text{in}}\geq m_{\text{out}}$. For our purposes, an interesting use of a collection autoencoder is to perform a ‘purification’ of the input samples in order to produce platonic [^13] By ‘platonic’ we mean ‘sharing the salient features with the majority of samples’ or quintessential. outputs. This is done by taking the size of the input space to be strictly larger than that of the output space, $m_{\text{in}}>m_{\text{out}}$, so that the autoencoder is forced to perform some averaging over the features of the input sample. In the ensuing sections we will frequently take advantage of this idea.
@@ 4.2 NCoder – $n$-point autoencoders
One common property of the autoencoders presented in 4.1 is that the encoder and decoder are trained simulataneously; that is the network’s optimiser updates the parameters of both the encoder and decoder functions in each iteration. A benefit of this approach is that the latent space is determined completely by the network, and is thus not free to be determined a-priori. At the same time, this can be disadvantageous in the sense that the user has little to no intuition as to how the autoencoder chooses to encode samples – autoencoders are ‘black-box’ systems.
In this note we introduce a novel autencoder construction with the intention of getting under the hood and understanding more about how complicated data can be efficiently encoded. We are interested in the case where the message space consists of images regarded as outputs from a lattice field theory as discussed in Section 3. Rather than leaving the construction of the latent space up to the model optimizer, we require the encoder to map collections of samples from the message space to a restricted set of sampled $n$-point functions, whilst the decoder analogously maps from the space of $n$-point functions back to (products of) the message space. To enforce the fixing of the latent space the $n$-point autoencoder will train the encoder an decoder functions separately.
@@
$y_{\text{in}}$ $\mapsto K^{\tilde{\mathcal{I}}}$ $K^{\tilde{\mathcal{I}}}$ $\mapsto z_{\text{out}}.$ (36)
Here $(y_{\text{in}})_{r}$ is a vector of flattened [^15] By ‘flattened’ we mean a vector which combines the sample index $j$ and the lattice site index $i$; this is merely introduced for computational convenience. input samples, and $K^{\tilde{\mathcal{I}}}$ is a vector of the associated $n$-point functions with respect to collections of lattice points contained in the index set $\tilde{\mathcal{I}}$. The flattening process simply vectorizes the training set such that
@@
We would like to stress the following interpretation for the autoencoder defined by (35). Firstly, the encoder is learning how to encode data in terms of their associated $n$-point correlation functions. Secondly, the decoder is learning how to reproduce data given the information contained in this correlation data. In this respect, it is useful to regard the $n$-point autoencoder as simulating a a statistical inference experiment. The latent layer, consisting of $n$-point correlation functions, should be interpreted as a set of (possibly) sufficient statistics characterizing the system of interest. The encoder function summarizes complicated data by reconstituting the information presented therein in terms of the sufficient statistics, while the decoder leverages said statistics in order to predict out of sample. In light of the discussion provided in Section 2.2, one may also wish to regard the autoencoder as solving the restricted moment/cumulant reconstruction problem. In this sense, the $n$-point autoencoder presents the further benefit of quantifying the explicit information contained in subsets of correlation functions.
@@ 4.2.1 Sampling Scheme
Let $T$ be a set of arbitrarily ordered image samples. The first step we will take to make the data more manageable is to train the network described in Section 4.2 in batches. Without loss of generality, $T$ can be decomposed into a union of batches $B^{\beta}$ such that
@@
Here, we have introduced the following notation: $B_{i}^{j;\beta}$ refers to the realized pixel output at the $i^{th}$ lattice site associated with the $j^{th}$ sample in the $\beta^{th}$ batch. Each batch consists of $N_{B}=|B|$ samples with $|B|$ chosen or $T$ slightly truncated such that $|T|\text{ mod }|B|=0$. By construction, each batch contains unique samples, i.e.
@@
Due to the typical size of image lattice being very large, the computation of $n$-point products over all lattice site combinations is computationally intensive. For example, in MNIST each image has
@@
3-point combinations – it is infeasible to consider them all for each batch. To overcome this difficulty, we need an approach which approximates the distribution of a carefully sampled subset of combinations. This subspace should be associated with an intentionally biased subset of points prioritizing the inclusion of points that concentrate around the subject of the image. As a result, highly-degenerate data is limited reducing the computational burden and better encoding the short-range information implicit in the images.
@@ 5.2 Decoder setup
The decoder network, whilst disparate, is taken to have a similar architecture to that of the encoder. Instead of mapping to a larger latent space, the decoder acts as a map from the space of $n$-point functions into the space of samples. In other words, the decoder attempts to reconstruct data which are consistent with the inserted summary statistics. As has been stressed in previous section, the message spaces of the encoder and decoder networks are different in our $N$-coder architecture. In particular, the encoder maps from a batch of samples, whilst the decoder maps to a single representative sample.
@@ 5.4 MNIST reconstructions
We are now prepared to apply our NCoder architecture to a specific data task, namely learning features of samples from the MNIST dataset. The network parameters are listed in Figure 9 in Appendix A. Figure 10 presents the training losses for the decoder of the neural networks encoding single point, single- and two-point functions, and single-, two-, and three-point functions.
In Figure 6 the reader can find reconstructed MNIST images trained on samples of the number seven. The reconstructed samples are organized from left to right by the highest order $n$-point correlation functions which are included in the latent layer of the NCoder network. By inspection, increasing the number of $n$-point functions improves the visual acuity of the reconstructed images. This is corroborated by validation measure in terms of the classifier accuracy, as can be seen in Table 5. Here, the validation accuracy is measured using a 2-way classifier technique described in detail in Appendix D. The stated accuracy is an average over 10 separate training runs of the classifier with random training sample ordering to ensure numerical stability.
@@ Acknowledgements
We wish to thank Jim Halverson and Anindita Maiti for introducing us to the Edgeworth expansion and explaining their work on the NN/FT correspondence at StringData 2023. We also wish to thank Jessica Howard for her suggestions about variations of the NCoder architecture and other particle physics oriented applications of autoencoders. Finally, we’d like to thank Edward Hirst for our many enlightening discussions and for all of his insightful comments. DSB and AGS acknowledge support from Pierre Andurand over the course of this research, and DSB is also supported by the Science and Technology Facilities Council (STFC) Consolidated Grant ST/T000686/1 “Amplitudes, strings & duality.” MSK is supported through the Physics department at the University of Illinois at Urbana-Champaign.
@@ Appendix B Cumulants in Quantum Field Theory and Statistical Mechanics
B.0.1 Statistical Mechanics
Consider the canonical ensemble, a system encompasing all states of a thermodynamic system which is in thermal equilibrium with a fixed-temperature heat bath. The canonical ensemble has an associated probability distribution of states governed by the thermodynmaic variable of temperature $T$. For convenience, we introduce the ‘thermodynamic beta’ $\beta=\frac{1}{k_{B}T}$, where $k_{B}$ is Boltzmann’s constant.
@@ B.0.1 Statistical Mechanics
$\frac{1}{Z[\beta]}\frac{\partial^{k}}{\partial t^{k}}Z[\beta-t]\bigg{\rvert}_{t=0}=(-1)^{k}\frac{Z^{(k)}[\beta]}{Z[\beta]}.$ (54)
@@
$\tilde{K}[\beta]:=-\ln Z[\beta]=-\beta^{-1}F$ (56)
meaning the cumulant generating functional is a multiple of the (Helmholtz) free energy [^21] Notice our definition of the Holmholtz energy differs by a factor of -1 to the usual definition. $F=\ln Z[\beta]$; the dimensionless free energy serves as a cumulant generating functional. The corresponding $k$-th cumulant is thus
@@
B.0.2 Quantum Field Theory

### v3

@@ 1 Introduction
To address our various questions we introduce in this note the NCoder – a neural network built around our aim for interpretability. The NCoder can be viewed as a deconstructed autoencoder in which the ‘latent layer’ is not built during the simultaneous optimization of the encoder/decoder but rather is fixed before training. Although this constraint limits the freedom of the network, it improves its ‘readability’. Motivated by the central role played by $n$-point functions in the analysis of field theories and general data science models, the NCoder uses the space of correlation functions as its prescribed latent layer. This set up renders the NCoder a potentially powerful tool for analyzing the importance of interpretable statistics in data generation and classification tasks. The paper is organized as follows: in Section 2 we review relevant concepts from probabiltiy and statistics such as the definition of moments and cumulants. In Section 2.2 we introduce the moment/cumulant reconstruction problem, and present a general solution in terms of the Edgeworth expansion which generalizes ideas from perturbation theory to arbitrary probability models. In Section 3 we briefly review how image data can be regarded as draws from a lattice field theory in preparation for the analysis of the latter sections. Section 4 presents the NCoder in detail, with a special focus on rendering the task of obtaining $n$-point functions feasible in light of the computational cost of the combinatorics involved. In Section 5 we use the NCoder to perform image generation tasks involving the MNIST [26] and Fashion MNIST [27] datasets. We conclude in Section 7 in which we discuss our findings, and suggest interesting directions for future work.
@@ 2.2 The reconstruction problem and the Edgeworth expansion
In what follows we will often work under the assumption that $p_{0}$ is a Gaussian distribution whose first and second cumulants match with those of $p$. In this case (20) takes the form
@@ 4.1 Review of autoencoders
In most cases $m_{\text{in}}\geq m_{\text{out}}$. For our purposes, an interesting use of a collection autoencoder is to perform a ‘purification’ of the input samples in order to produce platonic [^13] By ‘platonic’ we mean ‘sharing the salient features with the majority of samples’ or quintessential. outputs. To be more explicit, a collection autoencoder will extract salient features from the batch of input samples and produce a sample which is representative of the input distribution. For a sufficiently well-trained model, one may consider this as a type of averaging procedure, where the averaging occurs over the features of the distribution that the network learns during training. This is done by taking the size of the input space to be strictly larger than that of the output space, $m_{\text{in}}>m_{\text{out}}$, so that the autoencoder is forced to perform some averaging over the features of the input sample. One may implement this most simply in practice by modifying the loss of (32) such that $\mathcal{L}(y|\theta,\phi)\to\mathcal{N}\sum_{\beta}\mathcal{L}^{\beta}(y|\theta,\phi)$, where $\beta$ indexes each element in the collection. In the ensuing sections we will frequently take advantage of this idea.
@@ 4.2 NCoder – $n$-point autoencoders
One common property of the autoencoders presented in Section 4.1 is that the encoder and decoder are trained simulataneously; that is the network’s optimiser updates the parameters of both the encoder and decoder functions in each iteration. A benefit of this approach is that the latent space is determined completely by the network, and is thus not free to be determined a-priori. At the same time, this can be disadvantageous in the sense that the user has little to no intuition as to how the autoencoder chooses to encode samples – autoencoders are ‘black-box’ systems.
In this note we introduce a novel autencoder construction with the intention of getting under the hood and understanding more about how complicated data can be efficiently encoded. We are interested in the case where the message space consists of images regarded as outputs from a lattice field theory as discussed in Section 3. Rather than leaving the construction of the latent space up to the model optimizer, we require the encoder to map collections of samples from the message space to a restricted set of sampled $n$-point functions, whilst the decoder analogously maps from the space of $n$-point functions back to (products of) the message space. To enforce the fixing of the latent space the $n$-point autoencoder will train the encoder and decoder functions separately.
@@
$y_{\text{in}}$ $\mapsto K^{\tilde{\mathcal{I}}}$ $K^{\tilde{\mathcal{I}}}$ $\mapsto z_{\text{out}},$ (36)
where $\tilde{\mathcal{I}}$ is a subset of the full index set associated with all $n$-pt functions included in (27). Here $(y_{\text{in}})_{r}$ is a vector of flattened [^15] By ‘flattened’ we mean a vector which combines the sample index $j$ and the lattice site index $i$; this is merely introduced for computational convenience. input samples, and $K^{\tilde{\mathcal{I}}}$ is a vector of the associated $n$-point functions with respect to collections of lattice points contained in the index set $\tilde{\mathcal{I}}$. The flattening process simply vectorizes the training set such that
@@
We would like to stress the following interpretation for the autoencoder defined by (35). Firstly, the encoder is learning how to encode data in terms of their associated $n$-point correlation functions. Secondly, the decoder is learning how to reproduce data given the information contained in this correlation data. In this respect, it is useful to regard the $n$-point autoencoder as simulating a statistical inference experiment. The latent layer, consisting of $n$-point correlation functions, should be interpreted as a set of (possibly) sufficient statistics characterizing the system of interest. The encoder function summarizes complicated data by reconstituting the information presented therein in terms of the sufficient statistics, while the decoder leverages said statistics in order to predict out of sample. In light of the discussion provided in Section 2.2, one may also wish to regard the autoencoder as solving the restricted moment/cumulant reconstruction problem. In this sense, the $n$-point autoencoder presents the further benefit of quantifying the explicit information contained in subsets of correlation functions.
@@ 4.2.1 Sampling Scheme
Let $T$ be a set of arbitrarily ordered image samples. The first step we will take to make the data more manageable is to train the network described in Section 4.2 in batches. Without loss of generality, $T$ can be decomposed into a (disjoint) union of batches $B^{\beta}$ such that
@@
Here, we have introduced the following notation: $B_{i}^{\beta;j}$ refers to the realized pixel output at the $i^{th}$ lattice site associated with the $j^{th}$ sample in the $\beta^{th}$ batch. Each batch consists of $N_{B}=|B|$ samples with $|B|$ chosen or $T$ slightly truncated such that $|T|\text{ mod }|B|=0$. By construction, each batch contains unique samples, i.e.
@@
Due to the typical size of image lattices being very large, the computation of $n$-point products over all lattice site combinations is computationally intensive. For example, in MNIST each image has
@@
3-point combinations – it is infeasible to consider them all for each batch. To overcome this difficulty, we need an approach which approximates the distribution of a carefully sampled subset of combinations. This subspace should be associated with an intentionally biased subset of points prioritizing the inclusion of points that concentrate around the subject of the image. As a result, highly-degenerate data is limited, reducing the computational burden and better encoding the short-range information implicit in the images.
@@ 5.2 Decoder setup
The decoder network, whilst disparate, is taken to have a similar architecture to that of the encoder. Instead of mapping to a larger latent space, the decoder acts as a map from the space of $n$-point functions into the space of samples. In other words, the decoder attempts to reconstruct data which are consistent with the inserted summary statistics. As has been stressed in previous section, the message spaces of the encoder and decoder networks are different in our $N$-coder architecture.
@@ 5.4 MNIST reconstructions
We are now prepared to apply our NCoder architecture to a specific data task, namely learning features of samples from the MNIST dataset. In this demonstration, we consider a subset of the MNIST dataset consisting of a single class of samples: those depicting the digit ‘7’ [^20] One is free, of course, to choose any class of sample from the dataset. To be more concrete, the training batches $B^{\beta}$ are composed of drawn samples of sevens, each assumed to be IID. The network parameters are listed in Figure 12 in Appendix A. Figure 13 presents the training losses for the decoder of the neural networks encoding single point, single- and two-point functions, and single-, two-, and three-point functions.
In Figure 7 the reader can find reconstructed MNIST images trained on samples of the number seven. For reference, in Figure 5, we include an example batch of samples depicting the digit ‘7’ used for training. The reconstructed samples are organized from left to right by the highest order $n$-point correlation functions which are included in the latent layer of the NCoder network. By inspection, increasing the number of $n$-point functions improves the visual acuity of the reconstructed images. This is corroborated by the validation measure in terms of the classifier accuracy, as can be seen in Figure 6. Here, the validation accuracy is measured using a 2-way classifier technique described in detail in Appendix D. The stated accuracy is an average over 10 separate training runs of the classifier with random training sample ordering to ensure numerical stability.
Figure 5: An example of each samples of the digit ‘7’ in the MNIST dataset.
@@ 5.5 FashionMNIST reconstructions
6 Predicting the efficacy of the NCoder approach
When comparing the MNIST and Fashion MNIST results, we observe a stark difference in the efficacy of the NCoder reconstruction approach. In this section, we aim to understand this difference. The NCoder architecture is built upon the assumption that there is a hierarchy of importance amongst $n$-point correlation functions. In the implementation, one is free to cut off the architecture at a chosen $n$-point function with the expectation that higher $n$-point functions are less relevant for controlling the behaviour of the system under study. In the experiments reported upon in this note, we have truncated the NCoder at the $3$-point function, while also prioritizing the significance of the $1$- and $2$-point functions by limiting the number of $3$-point instances that are incorporated into the training sample [^21] In practice one must limit the number of $n-$ point functions for some value of $n$ as the number of samples to be included grows combinatorially.. This can be viewed as an assumption that the underlying data which is being modelled is essentially Gaussian (e.g. governed primarily by the $2$-point function) but with some small non-Gaussianities which the NCoder learns through the $3$-point function. From this point of view, the difference in the quality of the output between MNIST and Fashion MNIST reveals a difference in the underlying statistics of these datasets e.g. MNIST is well modelled by a GRP whereas Fashion MNIST should not be truncated in this way.
To test this hypothesis, we will now apply the NCoder architecture to modelling data in which we can control the ‘scale’ of deformations to an otherwise Gaussian system. Taking further inspiration from theoretical physics, we consider datasets which are comprised of samples of lattice configurations from the next-nearest-neighbour (NNN) Ising model. The classical Ising model in 2D is a model often used in statistical physics as a playground for studying phase transitions and critical phenomena. Spins are arranged on a 2D lattice, where each site has an associated ‘spin’ $\sigma_{i}=\pm 1$. In the presence of no external magnetic field, the model Hamiltonian is given by
$H=-J_{\text{NN}}\sum_{\braket{i,j}}\sigma_{i}\sigma_{j},$ (52)
where the notation $\braket{i,j}$ denotes a sum over all nearest neighbour pairs. It is typical to employ periodic boundary conditions such that each lattice site has the same number of neighbours.
The NNN Ising model is an extension of the classical Ising model with an additional term in the Hamiltonian incorporating next-nearest neighbour interactions. For equally spaced square lattices, the NNN interactions include the spins associated to the four nearest diagonal lattice sites (see Figure 8). For concreteness, the NNN Ising Hamiltonian is given by
$H=-J_{\text{NN}}\sum_{\braket{i,j}}\sigma_{i}\sigma_{j}-J_{\text{NNN}}\sum_{\braket{\braket{i,j}}}\sigma_{i}\sigma_{j},$ (53)
where $\braket{\braket{i,j}}$ represents all next-nearest-neighbour interactions. It is clear that sending $J_{\text{NNN}}\to 0$ recovers the nearest neighbour Ising model.
The NNN Ising model possesses two couplings, the ‘nearest neighbour’ coupling and the ‘next-nearest neighbour coupling’. In the limit in which $J_{\text{NNN}}<<J_{\text{NN}}$ we expect that the $2$-point function will be the dominant feature characterizing lattice configurations. As the next to nearest neighbor interaction grows in strength, higher point correlators will become comparable to the $2$-point function. So, if our hypothesis is correct, the efficacy of the NCoder model cutoff at the $2$-point function with minor corrections from a small number of $3$-point functions should diminish as a function of the coupling $J_{\text{NNN}}$.
Using Monte Carlo sampling, we consider a dataset constituting lattice spin configurations for a fixed set of parameters with only the next-nearest neighbour coupling varied. We perform this analysis in the low-temperature regime such that the effective temperature $T$ of the system is sufficiently low $T$ ($T=2$) and the model is in the ferromagnetic phase. Figure 9 shows the performance of the NCoder network as a function of increasing $J_{\text{NNN}}$. The vertical axis represents the KL divergence between input batches and output samples, averaged over 4000 samples. As expected, the KL-divergence increases monotonically.
This study provides clarity to the usefulness of the NCoder architecture, but also an indication of its possible pitfalls. Constructing an NCoder which depends upon some number of correlation functions provides a way of probing the statistical structure of a given sample. If the reconstruction is efficacious with some chosen cutoff – say up to $n^{*}$-point functions – we can argue that the data has an effective model in terms of $n$-point functions with $n\leq n^{*}$. On the other hand, however, if a dataset is not well approximated by a model which includes contributions from lower $n$-point functions we are forced to incorporate more and more $n$-point data into our reconstruction algorithm. Although including higher $n$-point functions is possible in principle, it increases the computational demand of the architecture. For example, even to implement the current instance of the NCoder which includes only a small number of $3$-point functions is computationally costly. Nevertheless, the NCoder remains a useful diagnostic tool (e.g. for determining the order or relevant interactions for a given dataset) even in cases where it is not computationally feasible to use it as a tool for full reconstruction.
Figure 8: Schematic of the nearest neighbouring spins (blue) and next-nearest neighbouring spins (red) for an arbitrary lattice site.
Figure 9: Relationship between KL divergence and the next nearest neighbour coupling $J_{\text{NNN}}$ for a set of 4000 samples over 5 averaged test networks.
@@ Acknowledgements
We wish to thank Jim Halverson and Anindita Maiti for introducing us to the Edgeworth expansion and explaining their work on the NN/FT correspondence at StringData 2023. We also wish to thank Jessica Howard for her suggestions about variations of the NCoder architecture and other particle physics oriented applications of autoencoders. Finally, we’d like to thank Edward Hirst for our many enlightening discussions, and insightful comments on the manuscript. DSB and AGS acknowledge support from Pierre Andurand over the course of this research, and DSB is also supported by the Science and Technology Facilities Council (STFC) Consolidated Grant ST/T000686/1 “Amplitudes, strings & duality.” MSK is supported through the Physics department at the University of Illinois at Urbana-Champaign.
@@ Appendix A Parameters and training losses for networks
Where stated, moving averages of window size 10 are used to smooth the noise in the training plots and improve readability.
@@ Appendix B Cumulants in Quantum Field Theory and Statistical Mechanics
B.1 Statistical Mechanics
Consider the canonical ensemble encompasing all states of a thermodynamic system which is in thermal equilibrium with a fixed-temperature heat bath. The canonical ensemble has an associated probability distribution of states governed by the thermodynmaic variable of temperature $T$. For convenience, we introduce the ‘thermodynamic beta’ $\beta=\frac{1}{k_{B}T}$, where $k_{B}$ is Boltzmann’s constant.
@@ B.1 Statistical Mechanics
$\frac{1}{Z[\beta]}\frac{\partial^{k}}{\partial t^{k}}Z[\beta-t]\bigg{\rvert}_{t=0}=(-1)^{k}\frac{Z^{(k)}[\beta]}{Z[\beta]},$ (56)
where the compact notation $Z^{(k)}[\beta]\equiv\frac{\partial^{k}}{\partial t^{k}}Z[\beta]$ has been introduced.
@@
$\tilde{K}[\beta]:=-\ln Z[\beta]=-\beta^{-1}F:=-\tilde{F}$ (58)
meaning the cumulant generating functional is a multiple of the (Helmholtz) free energy [^23] Notice our definition of the Helmholtz energy differs by a factor of -1 to the usual definition. $F=\ln Z[\beta]$; the dimensionless free energy $\tilde{F}=\beta F$ serves as a cumulant generating functional. The corresponding $k$-th cumulant is thus
@@
B.2 Quantum Field Theory